Operation Summary for September 2, 2011 - September 5, 2011 |
The Integrated Luminosity for the period from 8/29/11 to 9/5/11 was estimated at 58.92 pb-1 (inverse picobarns). NuMI reported receiving 8.08E18 protons on target during this same period. |
Friday September 2 The day shift began with Operations monitoring the Tevatron (TEV) store 9054 (71.69E30) and the Recycler stash (302.5E10), with the Antiproton Source (Pbar) stacking (8.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, at 8:56 AM, the I- Source vacuum valve closed, halting all beam; Linac personnel investigated and had beam back within 30 minutes. Operators reported that a TEV horizontal correction magnet power supply (T:HA22) was not following its reference; operators had its regulator changed out within 30 minutes. Operations terminated store 9054 at 12:29 PM. Operators put the TEV through a dry squeeze at 1:17 PM. Operations put the TEV into HEP shot setup at 1:41 PM and established store 9056 at 2:47 PM with an average initial luminosity (AIL) of 351.25E30. Operations established beam to the Meson FTBFT-1015 experiment at 2:55 PM. During the evening shift, Operators reported that the Computer Room temperatures had increased; FESS personnel investigated and found the air conditioning unit off and brought it back online, but the room temperatures did not recover. Operations disabled beam to T-1015 at 9 PM. Operators reported that the Booster RF system tripped off due to a CAMAC timer card problem; operators reset the card. Saturday September 3 The midnight shift began with Operations monitoring store 9056 (108.28E30) and the stash (223.8E10), with Pbar stacking (25.2E10), and with the NuMI and MiniBooNE taking beam. During the midnight shift, Operators reported that the Booster RF system tripped off again; operators replaced the CAMAC timer card. Operations terminated store 9056 at 7:45 AM and put the TEV into shot setup at 7:55 AM. During the day shift, Operations established store 9057 at 9:05 AM with an AIL of 361.0E30. FESS personnel discovered a faulty pump shaft had caused the Computer Room temperature problem; the temperatures recovered once they repaired the pump. Operations established beam to T-1015 at 10:45 AM. During the evening shift, Operations disabled beam to T-1015 at 7:10 PM. At 8:38 PM, operators reported that a MiniBooNE quadrupole magnet power supply (E:Q866D) had tripped off for the first of numerous times throughout the shift; it reset. Sunday September 4 The midnight shift began with Operations monitoring store 9057 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, operators changed out the switching power supply for E:Q866D. Operations terminated store 9057 at 12:30 AM and put the TEV into shot setup. At 1:15 AM, a TEV low beta magnet (DAU1) quenched due to an injection kicker (E48K) misfire. An EE Support engineer came in and pulsed the kicker, but it did not misfire again. Operations put the TEV back into shot setup at 5:10 AM and established store 9060 at 6 AM with an AIL of 363.0E30. During the day shift, at 9:51 AM, operators reported a Pbar beam permit trip due to a low target airflow indication. Operators called in a Mechanical Support person who replaced the blower motor. Pbar resumed stacking at 1:09 PM. During the evening shift, Operations terminated store 9060 at 11:58 PM. Monday September 5 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with the NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TEV into shot setup at 12:15 AM and established store 9061 at 1:15 AM with an AIL of 369.0E30. Operators reported that a Booster RF station (BRF7) tripped off numerous times throughout the shift. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 09/06/2011 email Fermilab |
| Security, Privacy, Legal | |